OpenStudy (anonymous):

Determine the density of NH3 gas at 435 K and 1.00 atm... I know that the ideal gas law is PV=nRT. I tried plugging in (1)V=(17)(0.082)(435). I got the volume (I think), but what do I get the density?

OpenStudy (anonymous):

Density is grams/volume (m/V) We can rewrite PV = nRT to find an exprssion that includes density. We know that moles * molecular weight = mass, or n * M = m so n = m/M Plug that into PV=nRT PV = nRT/M Now we notice that m/V is present if you rearrange: PM/(RT) =n/V = D (D for density) So PM/(RT) = D You can then plug in values to solve for density :)

OpenStudy (anonymous):

thank you very much! I think I just got a little bit confused about n and m/M. I'm still a little fuzzy on molarity.

OpenStudy (anonymous):

ohh okay. So M when we are talkin about PV=nRT and stoichiometry actually refers to molecular weight, not molarity. the units for molecular weight are g/mol That's why n * M = m mol *g/mol = g M, indeed, does mean molarity, when talking about concentration but that is a different area of chemistry :) Does that make any more sense now?

OpenStudy (anonymous):

ah yeah. That clears it up... so when I am calculating the molecular weight, those are the values I get when (for example) I convert 1 mol O2 to molecular weight (16*2)? Thank you :)

OpenStudy (anonymous):

yesiree! so in this case, you will calculate the moleular weight of NH3, so you do 14.01g/mol + 3(1.0079g)/mol which is approximately 17 grams/mol. This is M of NH3

OpenStudy (anonymous):

14.01 g/mol being the weight of the N and 1.0079 of H

OpenStudy (anonymous):

thank you so much! Now I understand my chemistry hw :p. You have a new fan :D

OpenStudy (anonymous):

awwww yay! I am glad that you understand your homework better now! that is my goal :)